OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Is this parallel lines in slope? If it is then the slope for the equation can be the same. So m=6

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

help